OpenStudy (anonymous):

Let's say we have intercepts of (9,0) and (0,-4), find the slope-intercept form of the line containing these two points.

OpenStudy (lgbasallote):

i prefer using the two-point-formula \[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\] \[\large \implies y - 0 = \frac{-4 - 0}{0 - 9} (x-9)\] \[\large \implies y = \frac{-4}{-9} (x-9)\] \[\large \implies y = \frac 49(x-9)\] \[\large \implies y = \frac 49 x - 4\] does that help?

OpenStudy (lgbasallote):

psst @warsop87 you there?

OpenStudy (anonymous):

Yea! Sorry thank you so much!!!

OpenStudy (lgbasallote):

welcome :D